# Calibrating Translation Decoding with Quality Estimation on LLMs

### v2

@@ Abstract
Neural machine translation (NMT) systems typically employ maximum a posteriori (MAP) decoding to select the highest-scoring translation from the distribution. However, recent evidence highlights the inadequacy of MAP decoding, often resulting in low-quality or even pathological hypotheses as the decoding objective is only weakly aligned with real-world translation quality. This paper proposes to directly calibrate hypothesis likelihood with translation quality from a distributional view by directly optimizing their Pearson correlation, thereby enhancing the effectiveness of translation decoding. With our method, translation with large language models (LLMs) improves substantially after limited training (2K instances per direction). This improvement is orthogonal to those achieved through supervised fine-tuning, leading to substantial gains across a broad range of metrics and human evaluations. This holds even when applied to top-performing translation-specialized LLMs fine-tuned on high-quality translation data, such as Tower, or when compared to recent preference optimization methods, like CPO. Moreover, the calibrated translation likelihood can directly serve as a strong proxy for translation quality, closely approximating or even surpassing some state-of-the-art translation quality estimation models, like CometKiwi. Lastly, our in-depth analysis demonstrates that calibration enhances the effectiveness of MAP decoding, thereby enabling greater efficiency in real-world deployment. The resulting state-of-the-art translation model, which covers 10 languages, along with the accompanying code and human evaluation data, has been released: https://github.com/moore3930/calibrating-llm-mt.
@@ 1 Introduction
The training of neural machine translation (NMT) has long been formulated as a maximum likelihood estimation (MLE) problem, and maximum a posteriori decoding (MAP) is a common decision rule, aiming to identify the highest-scoring translation. However, recent findings in NMT suggest that such a system has serious flaws. One particularly counterintuitive issue is the beam search curse (Koehn and Knowles, 2017; Murray and Chiang, 2018; Ott et al., 2018; Kumar and Sarawagi, 2019), where translation quality deteriorates as search approximations improve — higher-probability hypotheses are possibly worse translations.
Ott et al. (2018) explain this as a calibration issue — there is a generally low correlation between hypothesis likelihood and quality beyond a certain likelihood value. The flaws of poorly calibrated translation models can be categorized into two main aspects. First, their performance tends to be suboptimal, as better hypotheses may remain hidden within the distribution mass, inaccessible to MAP decoding (Ott et al., 2018). Second, likelihood, as a key measure of uncertainty, can serve as a valuable indicator of translation errors within the output at test time (Wang et al., 2020; Fomicheva et al., 2020); however, poorly calibrated models cannot effectively support this use.
Although these investigations were conducted on encoder-decoder architectures, we found that they also apply to translations elicited from LLMs. More specifically, we observed that (1) LLMs generally exhibit poor likelihood-quality calibration in translation, even when scaled up in capacity or continued pretrained with large amounts of multilingual data to enhance multilinguality; and (2) despite instruction fine-tuning with high-quality data leads to clear improvements in translation quality, as recommended by current best practices (Xu et al., 2023; Rei et al., 2024a), improvements in calibration are limited and sometimes barely observable.
Prior studies have tried to mitigate this miscalibration issue by introducing an additional optimization step during inference time, as shown in Figure 1 (top), known as quality-aware decoding (QAD) (Fernandes et al., 2022). These approaches typically involve generating multiple candidate translations through sampling, followed by reranking or voting using reference-free and/or reference-based machine translation metrics, such as Best-of-N (BoN) sampling (Rei et al., 2024a; Faria et al., 2024; Brown et al., 2024; Ichihara et al., 2025) and Minimum Bayes Risk (MBR) decoding (Kumar and Byrne, 2004; Freitag et al., 2022a). While effective, this line of methods inevitably incurs significantly higher latency. For example, the current top-performing system in the WMT24 competition (Kocmi et al., 2024a), Tower (Rei et al., 2024a), requires sampling 100 times for each prompt from a 70B LLM-based translation model; this is impractical, particularly for online environments.
This paper proposes a simple yet highly effective method to directly optimize the calibration during training, where we encourage the likelihood of translation hypotheses to align more closely with their quality distribution by directly optimizing Pearson correlations between them. More specifically, as shown in Figure 1 (bottom), given a translation prompt $x$, we sample multiple hypotheses from an NMT model during training time and compute both their likelihoods and quality scores, as measured by any external metric, such as COMET. We then define the loss as the negative Pearson correlation between the two sets of data points and minimize it directly using a standard gradient-based optimizer. While simple, our approach yields substantial performance improvements with limited training — using only 2K instances per language — even when applied to state-of-the-art LLM-based MT systems such as Tower. Moreover, due to clear calibration improvements, the resulting models’ likelihood can directly serve as a strong proxy for translation quality, even surpassing some state-of-the-art translation quality estimation (QE) models, like Comet-Kiwi (Rei et al., 2022a).
@@
Figure 1: The illustration of quality-aware decoding (top) and our calibration method (bottom). The former explores the decoding space by adding an extra step during test-time decoding, which involves multiple rounds of sampling followed by reranking or voting to select the highest-scoring output, incurring high costs. Our method focuses on training-time optimization, aiming to calibrate the likelihoods of hypotheses to their corresponding quality scores, enabling effective MAP decoding.
@@
  • We propose a simple yet effective method to mitigate the well-known miscalibration issue in translation during training time, where we calibrate translation decoding for quality from a holistic view by directly optimizing the Pearson correlation objective. With limited training, our models consistently outperform the strongest LLM-based MT systems – such as ALMA (Xu et al., 2023, 2024) and Tower (Rei et al., 2024a; Alves et al., 2024) series as well as the GPT-4o model — by a substantial margin across multiple automatic and human evaluation metrics. Extensive experiments demonstrate effectiveness over recent preference optimization methods, such as CPO (Xu et al., 2024), for translation.
@@ 2 Background
At a practical level, this study aims to optimize: (1) translation quality and (2) quality estimation, simultaneously in a single model. To this end, we briefly introduce related key foundational concepts.
@@
A few different themes investigate optimizing translation performance using signals from external metric models. We briefly summarize as follows:
(1) Test-time quality optimization. To alleviate the gap between decoding objective and translation quality, Minimum Bayes Risk (MBR) decoding (Kumar and Byrne, 2004; Freitag et al., 2022a; Yang et al., 2024) suggests basing decoding decisions on statistics gathered from the distribution holistically. As reference-free quality estimation (QE) progresses, Best-of-N sampling (Rei et al., 2024b; Faria et al., 2024; Brown et al., 2024; Ichihara et al., 2025) with a QE model becomes a straightforward strategy. However, both of these methods will result in a significant amount of additional decoding time. Moreover, the risk of metric bias increases, as translation improvements may stem from metric hacking rather than genuine quality enhancements (Skalse et al., 2022; Kovacs et al., 2024).
(2) Preference optimization. Recently, direct preference optimization (DPO) and its variants (Rafailov et al., 2023; Xu et al., 2024; Meng et al., 2024; Ethayarajh et al., 2024) have emerged as promising approaches for post-training LLMs, providing efficient alternatives to reinforcement learning from human feedback (RLHF). This line of methods often employs pairwise preferences under the Bradley-Terry framework (Bradley and Terry, 1952) to provide rewards, or a more general Plackett-Luce ranking framework (Plackett, 1975) when multiple ranked hypotheses are available — a few studies (Xu et al., 2024; Zhu et al., 2024) in translation also follow them. Additionally, some works (Guo et al., 2024; Lambert et al., 2024) employ these methods in an on-policy training paradigm for online LLMs alignments. Notably, regardless of employing pairwise/listwise data or on-/off-policy training, the objective is generally to maximize expected rewards. One key distinction of our method, as will be shown in §3, is that our training objective is to maximize correlation instead of expected rewards. Therefore, we follow prior research (Ott et al., 2018; Zhao et al., 2022) and use the term "calibration" in this paper to highlight the difference. More detailed discrepancies can be found in the next section.
@@ 3 Translation Calibration
Formally, given a parameterized auto-regressive language model $\theta$, and a translation instruction $x$, the log-likelihood of each translation hypothesis $y_{i}$ is denoted as $p_{\theta}(y_{i}|x)$. Meanwhile, the quality of this translation can be defined as $q(y_{i}|x)$ where $q$ represents any external quality evaluation model. When sampling hypotheses $y_{i}$ from $\theta$ conditioned on $x$, both $p_{\theta}(\cdot|x)$ and $q(\cdot|x)$ can be viewed as random variables defined over the output space.
Our goal is to calibrate the likelihoods of generated hypotheses with their quality — formally, to maximize the correlation between $p_{\theta}(\cdot|x)$ and $q(\cdot|x)$. Here, we bring the statistic, i.e., Pearson score $\rho(P,Q)$, to measure it. Pearson score $\rho(P,Q)$ measures the linear correlation between two variables $P$ and $Q$, which is defined as follows:
[…]
@@
where $\mathrm{cov}(P,Q)$ is the covariance between $\mathrm{P}$ and $\mathrm{Q}$, while $\sigma_{P}$ and $\sigma_{Q}$ are the corresponding standard deviations, and $\mu_{p}$ and $\mu_{q}$ are the corresponding means. This formulation computes the correlation by integrating the product of the centered values of $p$ and $q$ normalized by their standard deviations. Due to its scale-invariant properties and ability to measure trend consistency, the Pearson correlation coefficient is widely used in areas such as translation metric meta-evaluation.
@@
Given the intractably large decoding space, we employ sampling to approximate it. For each source sentence $x$, we first generate $k$ hypotheses $y_{i}$ ($i\in\{1,\dots,k\}$) by repeatedly prompting a large language model $\theta$ with nuclear sampling, and compute the corresponding $p_{\theta}(y_{i}|x)$. Also, we use a third-party metric model, like CometKiwi, to evaluate each hypothesis’s quality, i.e., $q(y_{i}|x)$.
By independently normalizing $p_{\theta}(\cdot|x)$ and $q(\cdot|x)$ to have zero mean and unit variance, we can define Pearson-based training loss for each source sentence $x$ as follows:
@@
$\mathcal{L}_{\text{pearson}}=-\frac{1}{k}\sum_{\begin{subarray}{c}i=1\\ y_{i}\sim p_{\theta}(y\mid x)\end{subarray}}^{k}\left(\frac{p_{\theta}(y_{i}\mid x)-\mu_{p}}{\sigma_{p}}\cdot\frac{q(y_{i}\mid x)-\mu_{q}}{\sigma_{q}}\right).$ (2)
@@
Note that we also introduce a supervised fine-tuning (SFT) term on the highest-scoring samples as a regularizer to ensure that the model’s likelihood distribution remains grounded in high-quality translations, as the Pearson objective alone only enforces correlation and does not constrain the absolute scale. The final loss for calibration is formulated as $\mathcal{L}_{\text{cal}}=\mathcal{L}_{\text{pearson}}+\mathcal{L}_{\text{sft}}$.
@@
An off-policy formulation can be obtained by trivially replacing the current model $\theta$ with an external model $\theta^{\ast}$ for sampling. Overall, by minimizing $\mathcal{L}_{\text{cal}}$ here, we encourage the Pearson score between $p$ and $q$ to be higher. In practice, we use a gradient-based optimizer, Adam, to optimize $\theta$ for this goal. Despite its simplicity, several important characteristics are captured in this formulation:
@@
  • It models hypothesis qualities from a holistic view, enabling the model to make finer-grained distinctions in translation quality within decoding space.
@@
  • Pearson’s correlation inherently applies normalization on a group of both likelihood and quality points. This normalization makes the objective invariant to scale and shift, thereby promoting stable and robust optimization across diverse input distributions.
  • The objective — Pearson’s score itself – is inherently shared with that of translation metric meta-evaluation, offering a unified perspective for both quality optimization and estimation.
Overall, the Pearson loss can be reduced to a mere dot product between two sets of normalized points. Despite this extreme simplicity, we show in the next section that it is highly effective in simultaneously optimizing translation quality and quality estimation within a single model.
@@ 4 Experimental Settings
4.1 Base Models
We base our experiments on two series of strong translation-specified LLMs, i.e., ALMA-Base (Xu et al., 2023) and Tower-Base (Alves et al., 2024), in both 7B and 13B variants. Both models achieve remarkable translation performance after post-training. For instance, the Tower series systems (Rei et al., 2024b) achieved state-of-the-art performance by winning all translation tasks in the WMT24 general translation tracks (Kocmi et al., 2024a), outperforming both Google Translate and GPT-4.
4.2 Data and Evaluation
Translation evaluation dataset. For fair comparison with the ALMA and Tower model series, we evaluate translation performance on the WMT22 and WMT24 datasets (Zerva et al., 2022; Kocmi et al., 2024a), respectively. Detailed dataset descriptions can be found in Appendix A. Following current best practice (Freitag et al., 2022b, 2023), all of the results in this paper are measured by widely used neural metrics, involving reference-based metric COMET, and the reference-free metrics XCOMET, CometKiwi-XL, and CometKiwi-XXL [^1] The corresponding metric model versions are Unbabel/wmt22-comet-da, Unbabel/XCOMET-XXL, Unbabel/wmt23-cometkiwi-da-xl, and Unbabel/wmt23-cometkiwi-da-xxl, respectively..
Metric meta-evaluation dataset. To assess how well translation likelihoods correlate with human judgments, we use the sentence-level quality estimation (QE) datasets during WMT-22 (Zerva et al., 2022), where sentence pairs are annotated in multiple ways to reflect translation quality, e.g., direct assessments (DA), post-edits (PE), and multidimensional quality metrics (MQM) (Freitag et al., 2021). In this paper, we use MQM scores as the ground-truth human annotations against which the metrics’ scores are evaluated, as MQM scores come from expert translators and are more reliable than the crowdsourced DA scores. Detailed description for the dataset can be found in Appendix A.
Calibration dataset. For the training set, we merge all English sentences from the Flores-200 dataset (Costa-Jussà et al., 2022) in dev and devtest splits and use them as the source, consisting of 2009 samples. For both on- and off-policy experiments, we use these sentences to construct translation prompts for each direction. The prompt templates can be found in Appendix E. For off-policy setting, we query gpt-4o-mini 16 times per prompt, employing nucleus sampling with a temperature of 1.0 and a top-p of 0.98. The resulting bitexts are evaluated using varying metrics to reflect corresponding quality scores. For on-policy experiments, all settings remain the same, except that a top-k value of 5 is used for each sampling step, and the model itself is queried directly.

### v3

@@ Abstract
Neural machine translation (NMT) systems typically employ maximum a posteriori (MAP) decoding to select the highest-scoring translation from the distribution. However, recent evidence highlights the inadequacy of MAP decoding, often resulting in low-quality or even pathological hypotheses as the decoding objective is only weakly aligned with real-world translation quality. This paper proposes to calibrate hypothesis likelihood with translation quality from a distributional view by directly optimizing their Pearson correlation, thereby enhancing decoding effectiveness. With our method, translation with large language models (LLMs) improves substantially after limited training (2K instances per direction). This improvement is orthogonal to those achieved through supervised fine-tuning, leading to substantial gains across a broad range of metrics and human evaluations. This holds even when applied to top-performing translation-specialized LLMs fine-tuned on high-quality translation data, such as Tower, or when compared to recent preference optimization methods, like CPO. Moreover, the calibrated translation likelihood can directly serve as a strong proxy for translation quality, closely approximating or even surpassing some state-of-the-art translation quality estimation models, like CometKiwi. Lastly, our in-depth analysis demonstrates that calibration enhances the effectiveness of MAP decoding, thereby enabling greater efficiency in real-world deployment. The resulting state-of-the-art translation model, which covers 10 languages, along with the accompanying code and human evaluation data, has been released: https://github.com/moore3930/calibrating-llm-mt.
@@ 1 Introduction
Figure 1: The illustration of quality-aware decoding (top) and our calibration method (bottom). The former explores the decoding space by adding an extra step during test-time decoding, which involves multiple rounds of sampling followed by reranking or voting to select the highest-scoring output. Our method focuses on training-time optimization, aiming to calibrate the likelihoods of hypotheses to their corresponding quality scores, enabling effective approximated MAP decoding.
The training of neural machine translation (NMT) has long been formulated as a maximum likelihood estimation (MLE) problem, and maximum a posteriori decoding (MAP) is a common decision rule, aiming to identify the highest-scoring translation. However, recent findings in NMT suggest that such a system has serious flaws. One particularly counterintuitive issue is the beam search curse (Koehn and Knowles, 2017; Murray and Chiang, 2018; Ott et al., 2018; Kumar and Sarawagi, 2019), where translation quality deteriorates as search approximations improve and higher-probability hypotheses are possibly worse translations.
Ott et al. (2018) explain this as a calibration issue. There is a generally low correlation between hypothesis likelihood and quality beyond a certain likelihood value. The flaws of poorly calibrated translation models can be categorized into two main aspects: First, their performance tends to be suboptimal, as better hypotheses may remain hidden within the distribution mass, inaccessible to MAP decoding (Ott et al., 2018). Second, likelihood, as a key measure of uncertainty, can serve as a valuable indicator of translation errors within the output at test time (Wang et al., 2020; Fomicheva et al., 2020); however, poorly calibrated models cannot effectively support this use.
Prior studies have tried to mitigate this miscalibration issue by introducing an additional optimization step during inference time, as shown in Figure 1 (top), known as quality-aware decoding (QAD) (Fernandes et al., 2022). These approaches typically involve generating multiple candidate translations through sampling, followed by reranking or voting using reference-free and/or reference-based machine translation metrics, such as Best-of-N (BoN) sampling (Rei et al., 2024a; Faria et al., 2024; Brown et al., 2024; Ichihara et al., 2025) and Minimum Bayes Risk (MBR) decoding (Kumar and Byrne, 2004; Freitag et al., 2022a). While effective, this line of methods inevitably incurs significantly higher latency. For example, the current top-performing system in the WMT24 competition (Kocmi et al., 2024a), Tower (Rei et al., 2024a), requires sampling 100 times for each prompt from a 70B LLM-based translation model, which is impractical, particularly for online environments.
This paper proposes a simple yet highly effective method to directly optimize the calibration during training, where we encourage the likelihood of translation hypotheses to align more closely with their quality by directly optimizing Pearson correlations between them. More specifically, as shown in Figure 1 (bottom), given a translation prompt $x$, we sample multiple hypotheses from an NMT model during training time and compute both their likelihoods and quality scores, as measured by any external metric, such as COMET. We then define the loss as the negative Pearson correlation between the two sets of data points and minimize it directly using a standard gradient-based optimizer. While simple, our approach yields substantial performance improvements with limited training—using only 2K instances per language—even when applied to state-of-the-art LLM-based MT systems such as Tower. Moreover, due to clear calibration improvements, the resulting models’ likelihood can directly serve as a strong proxy for translation quality, even surpassing some state-of-the-art translation quality estimation (QE) models, like Comet-Kiwi (Rei et al., 2022a).
@@
  • We introduce a simple yet effective method to mitigate the well-known miscalibration issue in translation during training time, where we calibrate translation decoding for quality from a holistic view by directly optimizing the Pearson correlation objective. With limited training, our models consistently outperform the strongest LLM-based MT systems, such as ALMA (Xu et al., 2023, 2024) and Tower (Rei et al., 2024a; Alves et al., 2024) series as well as the GPT-4o model, by a substantial margin across multiple automatic and human evaluation metrics. Extensive experiments demonstrate effectiveness over recent preference optimization methods, such as CPO (Xu et al., 2024), for translation.
@@ 2 Background
At a practical level, this study aims to optimize both translation quality as well as quality estimation, simultaneously in a single model. To this end, we briefly introduce the relevant foundational concepts.
@@
A few different themes investigate optimizing translation performance using signals from external metric models. We briefly summarize them as follows:
(1) Test-time quality optimization. To alleviate the gap between decoding objective and translation quality, Minimum Bayes Risk (MBR) decoding (Kumar and Byrne, 2004; Freitag et al., 2022a; Yang et al., 2024) suggests basing decoding decisions on statistics gathered from the distribution as a whole. As reference-free quality estimation (QE) progresses, Best-of-N sampling (Rei et al., 2024b; Faria et al., 2024; Brown et al., 2024; Ichihara et al., 2025) with a QE model becomes a straightforward strategy. However, both of these methods result in a significant amount of additional decoding time. Moreover, the risk of metric bias increases as translation improvements may stem from metric hacking rather than genuine quality enhancements (Skalse et al., 2022; Kovacs et al., 2024).
(2) Preference optimization. Recently, direct preference optimization (DPO) and its variants (Rafailov et al., 2023; Xu et al., 2024; Meng et al., 2024; Ethayarajh et al., 2024) have emerged as promising approaches for post-training LLMs, providing efficient alternatives to reinforcement learning from human feedback (RLHF). This line of methods often employs pairwise preferences under the Bradley-Terry framework (Bradley and Terry, 1952) to provide rewards, or a more general Plackett-Luce ranking framework (Plackett, 1975) when multiple ranked hypotheses are available and a few studies (Xu et al., 2024; Zhu et al., 2024) in translation also follow them. Additionally, some works (Guo et al., 2024; Lambert et al., 2024) employ these methods in an on-policy training paradigm for online LLM alignments. Notably, regardless of employing pairwise/listwise data or on-/off-policy training, the objective is generally to maximize expected rewards. A key distinction of our method, as detailed in Section 3, is that it optimizes correlation rather than expected rewards, forming the basis for unifying quality optimization and estimation. Therefore, we follow prior research (Ott et al., 2018; Zhao et al., 2022) and use the term “calibration” to highlight the difference. More detailed differences are discussed in the next section.
@@ 3 Translation Calibration
Formally, given a parameterized auto-regressive language model $p_{\theta}$ and a translation instruction $x$, the log-likelihood of a translation hypothesis $y_{i}$ is denoted as $z_{\theta}(y_{i}|x)=\log p_{\theta}(y_{i}|x)$. Meanwhile, the quality of this translation can be defined as $q(y_{i}|x)$ where $q$ represents any external quality evaluation model. When sampling hypotheses $y$ from $p_{\theta}$ conditioned on $x$, both $z_{\theta}(y|x)$ and $q(y|x)$ can be viewed as random variables defined over the output space. Our goal is to calibrate the likelihoods of generated hypotheses with their quality to maximize the correlation between $z_{\theta}(y|x)$ and $q(y|x)$.
Here, we use the statistic, i.e., Pearson correlation coefficient $\rho(a,b)$, to quantify the correlation. Let $a,b:\mathcal{Y}\to\mathbb{R}$ be two real-valued functions defined over a domain $\mathcal{Y}$. Their covariance with respect to a data distribution $p(y)$ is defined as
$\operatorname{cov}(a,b)=\mathbb{E}_{y\sim p}\left[\left(a(y)-\mathbb{E}_{y\sim p}[a(y)]\right)\left(b(y)-\mathbb{E}_{y\sim p}[b(y)]\right)\right].$ (1)
The corresponding Pearson score between $a$ and $b$ is given by
$\rho(a,b)=\frac{\operatorname{cov}(a,b)}{\sqrt{\operatorname{cov}(a,a)\operatorname{cov}(b,b)}}=\frac{\mathbb{E}_{y\sim p}\left[\left(a(y)-\mu_{a}\right)\left(b(y)-\mu_{b}\right)\right]}{\sigma_{a}\sigma_{b}},$ (2)
@@
where $\mu_{a}$, $\mu_{b}$ and $\sigma_{a}$, $\sigma_{b}$ denote expectations and standard deviations, respectively. This formulation computes the correlation by normalizing the expected product of their centered values. Due to its scale-invariance and ability to capture trend consistency, the Pearson correlation coefficient is widely used in translation metric meta-evaluation.
@@
In this study, we calculate and optimize $\rho$ with respect to the likelihood of hypotheses $z_{\theta}(y|x)$ and the quality score $q(y|x)$. Practically, given the intractably large decoding space, we employ Monte Carlo sampling for approximation. For each source sentence $x$, we generate $k$ hypotheses $y_{i}$ ($i\in\{1,\dots,k\}$) by repeatedly prompting a large language model $\theta$ with nucleus sampling, and compute the corresponding $z_{\theta}(y_{i}|x)$ and $q(y_{i}|x)$, and estimate the corresponding $\mu_{z}$, $\mu_{q}$ and $\sigma_{z}$, $\sigma_{q}$.
Notably, the standard definition of correlation assumes expectations under sampling from the full distribution, i.e., unweighted correlation. Here, however, we use nucleus sampling, which introduces a biased estimate by restricting sampling to high-probability regions. This approach is motivated by two factors: (1) uniform sampling over the output space is infeasible due to its vast size, and (2) following Ott et al. (2018), we focus on correlations among likely hypotheses. Accordingly, we define the Pearson-based loss using estimates under the nucleus-induced distribution $\tilde{p}$ as follows:
@@
$\mathcal{L}_{\text{pearson}}=-\frac{1}{k}\sum_{\begin{subarray}{c}i=1\\ y_{i}\sim\tilde{p}_{\theta}(\cdot|x)\end{subarray}}^{k}\left(\frac{z_{\theta}(y_{i}\mid x)-\mu_{z}}{\sigma_{z}}\cdot\frac{q(y_{i}\mid x)-\mu_{q}}{\sigma_{q}}\right).$ (3)
@@
We additionally introduce a supervised fine-tuning (SFT) term on the highest-scoring samples as a regularizer to ensure that the model’s likelihood distribution remains grounded in high-quality translations, since the Pearson objective alone enforces correlation but does not constrain the absolute scale. The final loss for calibration is formulated as $\mathcal{L}_{\text{cal}}=\mathcal{L}_{\text{pearson}}+\mathcal{L}_{\text{sft}}$.
@@
An off-policy formulation can be obtained by trivially replacing the current model $p_{\theta}$ with an external model $p_{\theta^{\ast}}$ for sampling. Overall, by minimizing $\mathcal{L}_{\text{cal}}$, we encourage the Pearson score between $z$ and $q$ to increase. In practice, we use a gradient-based optimizer, Adam, to optimize $\theta$ for this goal, with gradients propagated through $z_{\theta}$, $\mu_{z}$, and $\sigma_{z}$. Despite its simplicity, several important characteristics are captured in this formulation:
@@
  • It models hypothesis qualities from a holistic view, enabling the model to make finer-grained distinctions in translation quality within the decoding space.
@@
  • Pearson’s correlation inherently applies normalization to a group of both likelihood and quality points. This normalization makes the objective invariant to scale and shift, thereby promoting stable and robust optimization across diverse input distributions.
  • The objective, i.e., the Pearson’s score itself, is inherently shared with that of translation metric meta-evaluation, offering a unified perspective for both quality optimization and estimation. Meanwhile, unlike other statistics like Spearman’s or Kendall’s scores, Pearson’s coefficient is differentiable and thus suitable for gradient-based optimization frameworks.
Overall, the Pearson loss can be reduced to a mere dot product between two sets of normalized points, see Appendix C.1. Despite extreme simplicity, we show in the next sections that it is highly effective in simultaneously optimizing translation quality and quality estimation within a single model.
@@ 4 Experimental Settings
4.1 Base Models, Data, and Evaluation
Base Models. We base our experiments on two strong translation-specific LLMs, i.e., ALMA-Base (Xu et al., 2023) and Tower-Base (Alves et al., 2024), in their 7B and 13B variants. Both models achieve remarkable translation performance after post-training. For instance, the Tower series models (Rei et al., 2024b) achieved state-of-the-art results across all translation tasks in the WMT24 general translation tracks (Kocmi et al., 2024a), surpassing both Google Translate and GPT-4.
Translation evaluation dataset. For fair comparison with the ALMA and Tower model series, we evaluate translation performance on the WMT22 and WMT24 datasets (Zerva et al., 2022; Kocmi et al., 2024a), respectively. Detailed dataset descriptions can be found in Appendix D. Following current best practice (Freitag et al., 2022b, 2023), all results in this paper are measured by widely used neural metrics, involving reference-based metric COMET, and the reference-free metrics XCOMET, CometKiwi-XL, and CometKiwi-XXL [^1] The corresponding metric model versions are Unbabel/wmt22-comet-da, Unbabel/XCOMET-XXL, Unbabel/wmt23-cometkiwi-da-xl, and Unbabel/wmt23-cometkiwi-da-xxl, respectively..
Metric meta-evaluation dataset. To assess how well translation likelihoods correlate with human judgments, we use the sentence-level quality estimation (QE) datasets from WMT-22 (Zerva et al., 2022), where sentence pairs are annotated in multiple ways to reflect translation quality, e.g., direct assessments (DA), post-edits (PE), and multidimensional quality metrics (MQM) (Freitag et al., 2021). In this paper, we use MQM scores as ground-truth human annotations against which the metrics’ scores are evaluated, as MQM scores come from expert translators and are more reliable than the crowd-sourced DA scores. A detailed description of the dataset can be found in Appendix D.
Calibration dataset. For the training set, we merge all English sentences from the Flores-200 dataset (Costa-Jussà et al., 2022) in dev and devtest splits and use them as the source, consisting of 2,009 samples. For both on- and off-policy experiments, we use these sentences to construct translation prompts for each direction. The prompt templates can be found in Appendix E. For the off-policy setting, we query gpt-4o-mini 16 times per prompt, employing nucleus sampling with a temperature of 1.0 and a top-p of 0.98. The resulting bitexts are evaluated using varying metrics to reflect corresponding quality scores. For on-policy experiments, all settings remain the same, except that a top-k value of 5 is used for each sampling step, and the model itself is queried directly.